Two cars leave town headed north at the same time. Car A is moving at 60 mph, and Car B is moving at 80 mph. How long will it be until the cars are 55 miles apart?
Answer by mananth(16949)

same direction
subtract the speeds
80-60 = 20 mph
distance 55 miles
t= d/r
time = 55/20
=2.75 hours